OpenStudy (anonymous):

Anyone good with hyperbolas? Find the vertices and foci of the hyperbola with equation x squared over nine minus y squared over sixteen = 1

OpenStudy (anonymous):

\[\frac{ x^2 }{ 3^2 } -\frac{ y^2 }{ 4^2 } = 1\]

OpenStudy (anonymous):

Yes this was the problem we were working on last night. (: @saseal

OpenStudy (anonymous):